Training Data

Positive Training (22 cases)
Allah  Catholic  Christian  Dalai Lama  God
Jerry Falwell  Jesus  John the Baptist Mother Theresa Muhammad
Saint Jude  The Pope  Zeus  bible  church
crucifix  devout  holy  prayer  rabbi
religion  sacred
Negative Training (23 cases)
Abraham Lincoln Ben Franklin  Bill Clinton  Einstein  George Washington
Jimmy Carter  John Kennedy Michael Moore  atheist  dictionary
encyclopedia  evolution  helmet  internet  materialistic
minus  money  mouse  science  secular
seven  telephone  walking
Anchors  (6 dimensions)
evil follower history rational scripture

spirit

Testing Results

Positive tests  Negative tests
Positive altar, blessing, earth, shepherd
Predictions  communion, heaven,

sacrament, testament,

vatican

Negative angel Aristotle, Bertrand Russell,
Predictions  Greenspan, John,

Newton, Nietzsche, Plato, Socrates, air, bicycle, car, fire, five, man, monitor, water, whistle

Accuracy  24/27 = 88.89%